Match the annual percentage rate to each of these trade credit terms:

__ 1/5, NET 60 a. 12.29%

__ 1.5/10, NET 60 b. 10.11%

__ 1/15, NET 45 c. 11.12%

__ 1/10, NET 60 d. 6.70%

__ 1.5/5, NET 60 e. 7.37%
Answer #1

Cost of credit = (Discount %/(1-Discount %)) * 365/(Days taken-Discount period)

1:d

cost = 1%/99% * 365/(60-5) =6.7%

2: c

Cost = 1.5%/98.5% * 365/ (60-10) = 11.12%

3: a

Cost =1%/99%* 365/(45-15) =  12.29%

4: e

Cost = 1%/99% * 365/(60-10) = 7.37%

5: b

Cost = 1.5%/98.5%* 365/(60-5) = 10.11%